# Draw-In-Mind: Rebalancing Designer-Painter Roles in Unified Multimodal Models Benefits Image Editing

### v2

@@ Abstract
In recent years, integrating multimodal understanding and generation into a single unified model has emerged as a promising paradigm. While this approach achieves strong results in text-to-image (T2I) generation, it still struggles with precise image editing. We attribute this limitation to an imbalanced division of responsibilities. The understanding module primarily functions as a translator that encodes user instructions into semantic conditions, while the generation module must simultaneously act as designer and painter, inferring the original layout, identifying the target editing region, and rendering the new content. This imbalance is counterintuitive because the understanding module is typically trained with several times more data on complex reasoning tasks than the generation module. To address this issue, we introduce Draw-In-Mind (DIM), a dataset comprising two complementary subsets: (i) DIM-T2I, containing 14M long-context image–text pairs to enhance complex instruction comprehension; and (ii) DIM-Edit, consisting of 233K chain-of-thought imaginations generated by GPT-4o, serving as explicit design blueprints for image edits. We connect a frozen Qwen2.5-VL-3B (Bai et al., 2025) with a trainable SANA1.5-1.6B (Xie et al., 2025a) via a lightweight two-layer MLP, and train it on the proposed DIM dataset, resulting in DIM-4.6B-T2I/Edit. Despite its modest parameter scale, DIM-4.6B-Edit achieves SOTA or competitive performance on the ImgEdit and GEdit-Bench benchmarks, outperforming much larger models such as UniWorld-V1 (Lin et al., 2025) and Step1X-Edit (Liu et al., 2025). These findings demonstrate that explicitly assigning the design responsibility to the understanding module provides significant benefits for image editing. Our dataset and models are available at https://github.com/showlab/DIM.
@@ 1 Introduction
Motivated by this observation, we introduce Draw-In-Mind (DIM), a dataset consisting of two complementary subsets: (i) DIM-T2I that contains 14M long-context image–text pairs annotated across 21 dimensions by in-house models to lay the groundwork for complex chain-of-thought comprehension; and (ii) DIM-Edit that comprises 233K high-quality chain-of-thought imagination generated by GPT-4o from existing image editing data, enabling the model to interpret explicit design plans from an external designer. We then establish a simple baseline by concatenating a frozen MLLM, i.e., Qwen2.5-VL-3B, with a trainable DiT, i.e., SANA1.5-1.6B, via a two-layer MLP and train it end-to-end on both public data and the proposed DIM dataset, resulting in DIM-4.6B-T2I/Edit. During edit inference, we employ an arbitrary external designer, feeding its chain-of-thought imagination into the model to guide precise image edits. The framework and performance overview are illustrated in Figure 1. Despite its simplicity, DIM-4.6B-Edit matches or outperforms $5\times$ larger models such as Step1X-Edit (Liu et al., 2025) and UniWorld-V1 (Lin et al., 2025) on the ImgEdit benchmark. These results validate the effectiveness of the proposed DIM dataset and confirm the benefit of shifting the design responsibility from the generation module to the understanding module.
@@ 2.1 Existing Image Generation Datasets
T2I Datasets. Existing T2I datasets have provided many high-quality image–text pairs. They can be roughly grouped into three categories: (i) purely AI-generated data, e.g., JourneyDB (Sun et al., 2023) and MidJourney-V6 (CortexLM, 2025), which collect images from the MidJourney API, and HQ-Edit (Hui et al., 2024), which generates images via DALL-E 3; (ii) real-world data, e.g., COCO (Lin et al., 2014), which harvests images from Flickr and annotates them by human workers; and (iii) mixed data, e.g., InstructP2P (Brooks et al., 2023), which sources real images from LAION-Aesthetics and produces edited variants via Prompt2Prompt (Hertz et al., 2022). Although these datasets deliver high perceptual quality, their prompts are typically short, limiting their utility for complex chain-of-thought reasoning in the editing stage. To ensure broad concept coverage, we opt for harvesting real-world images and annotate them with our in-house models from 21 dimensions, yielding 14M long-context image–text pairs, namely DIM-T2I, that form a robust foundation for complex CoT-guided editing.
@@ 2.2 Unified Models for Image Generation
Image Editing Models. When it comes to image editing, the challenge becomes significantly harder, as neither the latest integrative models (Lin et al., 2025) nor connector-based ones (Liu et al., 2025) achieve satisfactory performance on mainstream benchmarks such as ImgEdit and GEdit-Bench compared to proprietary models like GPT-4o-Image, even when employing large-scale understanding and generation models such as Qwen2.5-VL-7B (Bai et al., 2025) and FLUX.1-dev (Labs, 2024). This suggests that simply scaling model size is not an effective strategy for improving image editing capability. In this work, we take a different approach by addressing the problem from a perspective of imbalanced division of responsibilities. We propose DIM-4.6B-Edit, which leverages an external designer to create blueprints in a CoT manner in the textual space before editing. Despite having only 1.6B generative parameters, our model achieves SOTA editing performance, highlighting the effectiveness of shifting the design responsibility to the understanding module.
@@ 3.1.1 DIM-T2I
Current image generation datasets perform well in terms of prompt–image alignment and image perceptual quality. However, the prompts in existing datasets are typically short and simple, as shown in Table 1. While these prompts accurately capture the semantics of the target image, they fall short in fostering long-context comprehension, which is an essential foundation for complex CoT–guided image editing. To bridge this gap, we collect 14M images with resolutions higher than $512\times 512$ from the web and annotate them using internal models across 21 dimensions listed in Appendix B, resulting in DIM-T2I. Our produced annotations are long and detailed, thoroughly covering all dimensions. As shown in Table 1, the average prompt length is at least twice that of existing corpora, effectively establishing a strong basis for complex CoT-guided image editing.
@@
It is worth noting that different from existing datasets that typically apply aesthetic filtering (Fei et al., 2024; Brooks et al., 2023) or ensure perceptual quality by generating images directly with AI (Sun et al., 2023; CortexLM, 2025), our DIM-T2I is collected directly from the real world without any post-filtering, ensuring broad coverage of world knowledge. DIM-T2I demonstrates that even without complex perceptual-level filtering on the visual side, lower-quality images paired with deliberately curated long-context prompts can still serve as an effective corpus for T2I generation, evidenced by the SOTA performance of DIM-4.6B-T2I trained on this dataset.
@@ 3.1.2 DIM-Edit
Finally, we provide the optimized prompts, along with the source image to GPT-4o and instruct it to produce a four-step CoT imagination that emulates human editing behavior. For the sake of accuracy, we also provide the target image to it for reference. The target of each CoT step is as follows: (i) Global Layout Perception: identify and describe all key objects and their positions in the source image. (ii) Local Object Perception: describe the appearance of each object or background element in the source image, including shape, color, texture, and state. (iii) Edit Area Localization: specify which objects or regions will be modified, based on the refined instruction. (iv) Edited Image Imagination: describe the expected appearance of the edited image, with an emphasis on the modified areas. As shown in Table 1 and Figure 2, the resulting CoT imagination is not only ultra-long but also highly accurate, effectively removing the design responsibility from the generation module and thereby significantly enhancing the efficiency of image editing learning. A quality assessment of the CoTs can be found in Appendix D.
@@
For the base T2I model, we start by establishing a simple baseline, similar to MetaQuery (Pan et al., 2025), to preserve state-of-the-art understanding capability. We select Qwen2.5-VL-3B (Bai et al., 2025) as the MLLM and SANA1.5-1.6B (Xie et al., 2025a) as the diffusion decoder for their modest size. Unlike MetaQuery, which employs a large 24-layer transformer with 1.6B parameters as a connector between the MLLM and the diffusion decoder, we adopt a much simpler design, i.e., a two-layer MLP, to directly project multimodal tokens into the generation space. We refer to this model as DIM-4.6B-T2I, illustrated in Figure 1. We train DIM-4.6B-T2I on a mixture of the proposed DIM-T2I dataset and an additional 6.9M image–text pairs from MidJourney-V6 (CortexLM, 2025), COCO (Lin et al., 2014), InstructP2P (Brooks et al., 2023), JourneyDB (Sun et al., 2023), HQ-Edit (Hui et al., 2024), and Dimba (Fei et al., 2024). During training, Qwen2.5-VL-3B remains frozen, and we finetune only the parameters of the connector and SANA1.5-1.6B. Notably, we exclude distillation datasets such as BLIP3-o-60K (Chen et al., 2025a) to avoid any risk of data leakage (Wu et al., 2025b) in the following evaluation. We adopt vanilla flow matching as the sole objective, avoiding parameter-tuning tricks to highlight data effectiveness and maintain simplicity.
Thanks to the rich world knowledge and high-quality long-context prompts in the proposed DIM-T2I dataset, the trained DIM-4.6B-T2I model provides a strong foundation for complex CoT comprehension. We then adopt a two-stage training strategy to adapt it for the editing task. In the first stage, we train DIM-4.6B-T2I on the UltraEdit (Zhao et al., 2024) dataset to develop basic editing capability. Following InstructP2P (Brooks et al., 2023), we concatenate the source image with noise along the channel dimension, as illustrated in Figure 1. In the second stage, we fine-tune the stage-one model exclusively on the proposed DIM-Edit dataset, resulting in DIM-4.6B-Edit. During inference, we employ an external designer, e.g., GPT-4o, to prepare a blueprint in the same format as DIM-Edit, except without access to the target image, ensuring alignment with real usage scenarios.
@@ 4.1 Experimental Setup
Although the primary focus of this paper is image editing, we also evaluate DIM-4.6B-T2I on T2I generation to verify the effectiveness of the proposed DIM-T2I. Specifically, we assess semantic alignment using GenEval (Ghosh et al., 2023) scores and perceptual quality using the FID on the MJHQ-30K (Li et al., 2024). Following MetaQuery (Pan et al., 2025) and Emu3 (Pan et al., 2025), we leverage an LLM to rewrite the original prompt for GenEval evaluation. For image editing, we report scores on the recently proposed ImgEdit (Lin et al., 2025) and GEdit-Bench-EN (Liu et al., 2025) benchmarks, using GPT-4.1 for evaluation to ensure fair comparison with existing results.
@@ 4.2.2 Image Editing
We further demonstrate the capability of DIM-4.6B-Edit through intuitive visual comparisons of editing results on four AI-generated out-of-domain images in Figure 3. Janus-4o exhibits severe distortions despite being trained on GPT-4o-generated edit pairs, while Step1X-Edit produces less natural edits (rows 2–4) and fails in complex scenarios such as row 1, which involves manipulating multiple objects. In contrast, DIM-4.6B-Edit successfully follows the instructions to produce natural and consistent edited images. Please refer to Appendix A for more visualizations.
@@ 4.3 Ablation Study
Designer Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
@@
Generalizability. Although our proposed DIM-Edit is annotated with GPT-4o to ensure data quality, we show that the resulting DIM-4.6B-Edit is compatible with various external designers, as reported in Table 5. In the first row, we remove the designer and directly use the raw prompt from ImgEdit. Even under this setting, DIM-4.6B-Edit achieves performance comparable to frontier models such as BAGEL, demonstrating that high-quality CoT annotations help strengthen basic editing by mitigating prompt–edit misalignment. We then replace GPT-4o with five mainstream MLLMs as external designers, i.e., Qwen2.5-VL-3B/7B (Bai et al., 2025), MiMo-VL-7B (Xiaomi, 2025), InternVL3.5-8B (Bai et al., 2025), and GLM-4.1V-9B (Hong et al., 2025). All of them deliver strong results compared to previous state-of-the-art models in Table 3, highlighting the robustness of DIM-4.6B-Edit and the generalizability of our DIM framework. Furthermore, models equipped with external designers significantly outperform the raw-prompt setting, confirming that CoT imagination effectively reduces the burden on the generation modules and enhances overall editing quality.

### v3

# arXiv:2509.01986v3[cs.CV] 27 Feb 2026

@@ Abstract
In recent years, integrating multimodal understanding and generation into a single unified model has emerged as a promising paradigm. While this approach achieves strong results in text-to-image (T2I) generation, it still struggles with precise image editing. We attribute this limitation to an imbalanced division of responsibilities. The understanding module primarily functions as a translator that encodes user instructions into semantic conditions, while the generation module must simultaneously act as designer and painter, inferring the original layout, identifying the target editing region, and rendering the new content. This imbalance is counterintuitive because the understanding module is typically trained with several times more data on complex reasoning tasks than the generation module. To address this issue, we introduce Draw-In-Mind (DIM), a dataset comprising two complementary subsets: (i) DIM-T2I, containing 14M long-context image-text pairs to enhance complex instruction comprehension; and (ii) DIM-Edit, consisting of 233K chain-of-thought imaginations generated by GPT-4o, serving as explicit design blueprints for image edits. We connect a frozen Qwen2.5-VL-3B (Bai et al., 2025) with a trainable SANA1.5-1.6B (Xie et al., 2025a) via a lightweight two-layer MLP, and train it on the proposed DIM dataset, resulting in DIM-4.6B-T2I/Edit. Despite its modest parameter scale, DIM-4.6B-Edit achieves SOTA or competitive performance on the ImgEdit and GEdit-Bench benchmarks, outperforming much larger models such as UniWorld-V1 (Lin et al., 2025) and Step1X-Edit (Liu et al., 2025). These findings demonstrate that explicitly assigning the design responsibility to the understanding module provides significant benefits for image editing. Our dataset and models are available at https://github.com/showlab/DIM.
@@ 1 Introduction
Motivated by this observation, we introduce Draw-In-Mind (DIM), a dataset consisting of two complementary subsets: (i) DIM-T2I that contains 14M long-context image-text pairs annotated across 21 dimensions by in-house models to lay the groundwork for complex chain-of-thought comprehension; and (ii) DIM-Edit that comprises 233K high-quality chain-of-thought imagination generated by GPT-4o from existing image editing data, enabling the model to interpret explicit design plans from an external designer. We then establish a simple baseline by concatenating a frozen MLLM, i.e., Qwen2.5-VL-3B, with a trainable DiT, i.e., SANA1.5-1.6B, via a two-layer MLP and train it end-to-end on both public data and the proposed DIM dataset, resulting in DIM-4.6B-T2I/Edit. During edit inference, we employ an arbitrary external designer, feeding its chain-of-thought imagination into the model to guide precise image edits. The framework and performance overview are illustrated in Figure 1. Despite its simplicity, DIM-4.6B-Edit matches or outperforms $5\times$ larger models such as Step1X-Edit (Liu et al., 2025) and UniWorld-V1 (Lin et al., 2025) on the ImgEdit benchmark. These results validate the effectiveness of the proposed DIM dataset and confirm the benefit of shifting the design responsibility from the generation module to the understanding module.
@@ 2.1 Existing Image Generation Datasets
T2I Datasets. Existing T2I datasets have provided many high-quality image-text pairs. They can be roughly grouped into three categories: (i) purely AI-generated data, e.g., JourneyDB (Sun et al., 2023) and MidJourney-V6 (CortexLM, 2025), which collect images from the MidJourney API, and HQ-Edit (Hui et al., 2024), which generates images via DALL-E 3; (ii) real-world data, e.g., COCO (Lin et al., 2014), which harvests images from Flickr and annotates them by human workers; and (iii) mixed data, e.g., InstructP2P (Brooks et al., 2023), which sources real images from LAION-Aesthetics and produces edited variants via Prompt2Prompt (Hertz et al., 2022). Although these datasets deliver high perceptual quality, their prompts are typically short, limiting their utility for complex chain-of-thought reasoning in the editing stage. To ensure broad concept coverage, we opt for harvesting real-world images and annotate them with our in-house models from 21 dimensions, yielding 14M long-context image-text pairs, namely DIM-T2I, that form a robust foundation for complex CoT-guided editing.
@@ 2.2 Unified Models for Image Generation
Image Editing Models. When it comes to image editing, the challenge becomes significantly harder, as neither the latest integrative models (Lin et al., 2025) nor connector-based ones (Liu et al., 2025) achieve satisfactory performance on mainstream benchmarks such as ImgEdit and GEdit-Bench compared to proprietary models like GPT-4o-Image, even when employing large-scale understanding and generation models such as Qwen2.5-VL-7B (Bai et al., 2025) and FLUX.1-dev (Labs, 2024a). This suggests that simply scaling model size is not an effective strategy for improving image editing capability. In this work, we take a different approach by addressing the problem from a perspective of imbalanced division of responsibilities. We propose DIM-4.6B-Edit, which leverages an external designer to create blueprints in a CoT manner in the textual space before editing. Despite having only 1.6B generative parameters, our model achieves SOTA editing performance, highlighting the effectiveness of shifting the design responsibility to the understanding module.
@@ 3.1.1 DIM-T2I
There are typically two strategies to train an editing model, i.e., (i) learning drawing first (T2I), followed by adaptation for editing, and (ii) directly learning editing. We observe that the vast majority of image editing models are built upon established T2I foundations (Brooks et al., 2023; Zhao et al., 2024; Liu et al., 2025). This aligns with the first strategy and represents a robust technical route that benefits from curriculum learning. Consequently, we elected to achieve basic T2I ability and subsequently fine-tune the base model for the more challenging editing task.
However, we observed that despite the current T2I datasets performing well in terms of prompt-image alignment and image perceptual quality, the prompts in existing datasets are typically short and simple, as shown in Table 1. While these prompts accurately capture the semantics of the target image, they fall short in fostering long-context comprehension, which is an essential foundation for complex CoT-guided image editing. To bridge this gap, we collect 14M images with resolutions higher than $512\times 512$ from the web. We believe that the aspects emphasized in widely recognized understanding datasets and benchmarks effectively capture the most frequent interactions between humans and objects in the real world. Therefore, we conduct a thorough literature review and an empirical analysis of existing understanding datasets and benchmarks, and finally derive 21 diverse dimensions and use internal models to generate long and detailed annotations, thoroughly covering all dimensions, resulting in DIM-T2I. As shown in Table 1, its average prompt length is at least twice that of existing corpora, effectively establishing a strong basis for complex CoT-guided image editing. The dimension-specific prompts and referred datasets/benchmarks are listed in Appendix E.
@@ 3.1.2 DIM-Edit
Finally, we provide the optimized prompts, along with the source image to GPT-4o and instruct it to produce a four-step CoT imagination that emulates human editing behavior. For the sake of accuracy, we also provide the target image to it for reference. The target of each CoT step is as follows: (i) Global Layout Perception: identify and describe all key objects and their positions in the source image. (ii) Local Object Perception: describe the appearance of each object or background element in the source image, including shape, color, texture, and state. (iii) Edit Area Localization: specify which objects or regions will be modified, based on the refined instruction. (iv) Edited Image Imagination: describe the expected appearance of the edited image, with an emphasis on the modified areas. As shown in Table 1 and Figure 2, the resulting CoT imagination is not only ultra-long but also highly accurate, effectively removing the design responsibility from the generation module and thereby significantly enhancing the efficiency of image editing learning. A quality assessment of the CoTs involving both MLLMs and human verification can be found in Appendix D.
@@
For the base T2I model, we start by establishing a simple baseline, similar to MetaQuery (Pan et al., 2025), to preserve state-of-the-art understanding capability. We select Qwen2.5-VL-3B (Bai et al., 2025) as the MLLM and SANA1.5-1.6B (Xie et al., 2025a) as the diffusion decoder for their modest size. Unlike MetaQuery, which employs a large 24-layer transformer with 1.6B parameters as a connector between the MLLM and the diffusion decoder, we adopt a much simpler design, i.e., a two-layer MLP, to directly project multimodal tokens into the generation space. We refer to this model as DIM-4.6B-T2I, illustrated in Figure 1. We train DIM-4.6B-T2I on a mixture of the proposed DIM-T2I dataset and an additional 6.9M image-text pairs from MidJourney-V6 (CortexLM, 2025), COCO (Lin et al., 2014), InstructP2P (Brooks et al., 2023), JourneyDB (Sun et al., 2023), HQ-Edit (Hui et al., 2024), and Dimba (Fei et al., 2024). During training, Qwen2.5-VL-3B remains frozen, and we finetune only the parameters of the connector and SANA1.5-1.6B. Notably, distillation datasets like BLIP3-o-60K (Chen et al., 2025a) explicitly curate data to align with the structural patterns of benchmarks like GenEval, we exclude them to avoid any risk of data leakage (Wu et al., 2025b) or benchmark hacking in the evaluation to justify the contribution of our DIM data. We adopt vanilla flow matching as the sole objective, avoiding parameter-tuning tricks to highlight data effectiveness and maintain simplicity.
Thanks to the rich world knowledge and high-quality long-context prompts in DIM-T2I, the trained DIM-4.6B-T2I model provides a strong foundation for complex CoT comprehension. We then adopt a two-stage training strategy to adapt it for the editing task. In the first stage, we initialize the editing model from DIM-4.6B-T2I and fine-tune it on the UltraEdit (Zhao et al., 2024) dataset to develop basic editing capability. Following InstructP2P (Brooks et al., 2023), we concatenate the source image with noise along the channel dimension, as illustrated in Figure 1. In the second stage, we fine-tune the stage-one model exclusively on the proposed DIM-Edit dataset, resulting in DIM-4.6B-Edit. During inference, we employ an external designer to prepare a blueprint in the same format as DIM-Edit, except without access to the target image, ensuring alignment with real usage scenarios.
@@ 4.1 Experimental Setup
Although the primary focus of this paper is image editing, we evaluate DIM-4.6B-T2I on T2I benchmarks to verify the effectiveness of DIM-T2I. We report the GenEval (Ghosh et al., 2023) scores and MJHQ-30K (Li et al., 2024b) FID. Following MetaQuery (Pan et al., 2025) and Emu3 (Pan et al., 2025), we test LLM-rewritten prompts for GenEval evaluation. For image editing, we report scores on the recently proposed ImgEdit (Lin et al., 2025) and GEdit-Bench-EN (Liu et al., 2025) benchmarks, using GPT-4.1 for evaluation to ensure fair comparison with existing results. We also report results on MagicBrush (Zhang et al., 2023) to show the performance on automated metrics.
@@ 4.2.2 Image Editing
We further demonstrate the capability of DIM-4.6B-Edit through intuitive visual comparisons of editing results on four AI-generated out-of-domain images in Figure 3. Janus-4o exhibits severe distortions despite being trained on GPT-4o-generated edit pairs, while Step1X-Edit produces less natural edits (rows 2-4) and fails in complex scenarios such as row 1, which involves manipulating multiple objects. In contrast, DIM-4.6B-Edit successfully follows the instructions to produce natural and consistent edited images. Please refer to Appendix A for more visualizations.
@@
Table 5: The MagicBrush test set performance. Metrics are calculated between human-edited groundtruth and AI-generated edits. [Uncaptioned image] and [Uncaptioned image] denote the 1st and 2nd best model, respectively.
[…]
InstructP2P (Brooks et al., 2023) 0.9B 0.114 0.851 0.744
MagicBrus (Zhang et al., 2023) 0.9B 0.074 0.908 0.847
UltraEdit (Zhao et al., 2024) 2.0B 0.066 0.904 0.852
FluxEdit (Paul, 2025) 12.0B 0.114 0.779 0.663
FLUX.1 Fill (Labs, 2024b) 12.0B 0.192 0.795 0.669
RF-Solver Edit (Wang et al., 2024a) 12.0B 0.112 0.766 0.675
ACE++ (Mao et al., 2025) 12.0B 0.195 0.741 0.591
ICEdit (Zhang et al., 2025) 12.0B 0.060[Uncaptioned image] 0.928[Uncaptioned image] 0.853[Uncaptioned image]
DIM-4.6B-Edit 1.6B 0.065[Uncaptioned image] 0.928[Uncaptioned image] 0.882[Uncaptioned image]
[…]
@@ 4.3 Ablation Study
External Designer Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
@@
Generalizability to External Designers. Although our proposed DIM-Edit is annotated with GPT-4o, we show that the resulting DIM-4.6B-Edit is compatible with various external designers, as reported in Table 6. In the first row, we remove the designer and directly use the raw prompt from ImgEdit. Even under this setting, DIM-4.6B-Edit achieves performance comparable to frontier models such as BAGEL, demonstrating that high-quality CoT annotations help strengthen basic editing by mitigating prompt–edit misalignment. We then replace GPT-4o with four mainstream MLLMs as external designers, i.e., Qwen2.5-VL-7B (Bai et al., 2025), MiMo-VL-7B (Xiaomi, 2025), InternVL3.5-8B (Bai et al., 2025), and GLM-4.1V-9B (Hong et al., 2025). All of them deliver strong results compared to previous state-of-the-art models in Table 3, highlighting the robustness of DIM-4.6B-Edit and the generalizability of our DIM framework. Furthermore, models equipped with external designers significantly outperform the raw-prompt setting, confirming that CoT imagination effectively reduces the burden on the generation modules and enhances overall editing quality.
Table 7: The ImgEdit performance w.r.t. the internal Qwen2.5-VL-3B designer.
Internal Designer Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
- 3.53 3.23 2.01 3.49 1.47 3.42 4.79 2.35 3.64 3.10
Qwen2.5-VL-3B[Uncaptioned image] 3.80 3.24 2.03 3.89 3.21 3.52 4.92 2.71 4.05 3.49
Qwen2.5-VL-3B[Uncaptioned image] 3.96 3.36 2.25 3.98 3.31 3.81 4.95 2.83 4.02 3.61
GPT-4o 4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
@@
Integrated End-to-End Evaluation. To exclude potential influence from external designers, we establish a “self-play” configuration. In this setup, CoT embeddings generated by the internal MLLM (Qwen2.5-VL-3B) are directly fed into the painter to execute edits, effectively eliminating the need for the external inference round. The result (Table 7 2nd row) shows that this “self-play” model achieves SOTA performance, validating the effectiveness of the Draw-In-Mind paradigm and the high quality of the DIM-Edit data. We further investigate whether the CoT blueprints in DIM-Edit can serve as a corpus to bridge the gap between open-source and closed-source designers. To this end, we perform lightweight fine-tuning on Qwen2.5-VL-3B and subsequently feed its blueprints into DIM-4.6B-Edit. The results (Table 7 3rd row) demonstrate that fine-tuning from DIM-Edit’s CoTs can effectively mitigate the performance disparity with the proprietary models like GPT-4o.
@@
Table 10: The ImgEdit performance of models initialized from scratch/DIM-4.6B-T2I.
Initialization Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
Scratch 2.70 2.56 1.93 2.23 2.47 2.82 4.68 2.38 2.15 2.66
DIM-4.6B-T2I 4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
Necessity of DIM-T2I. Our CoT-guided editing requires robust comprehension capabilities. We posit that T2I generation is simpler than editing and is better suited for fostering this capability. Rather than simultaneously tackling two challenging objectives, i.e., complex instruction comprehension and image editing, we chose to establish strong instruction comprehension first in the T2I stage. To justify our assumption, we trained a model exclusively on DIM-Edit to test the feasibility of simultaneously achieving modality alignment, complex instruction comprehension, and editing capabilities in a single stage. As evident from the Table 10, the model trained from scratch significantly underperforms the version initialized with DIM-4.6B-T2I. This performance gap empirically validates the necessity of DIM-T2I as a foundational cornerstone for the Draw-In-Mind paradigm.